Determine the area under the standard normal curve that lies between the following values.

zequals=1

and

zequals=2

2)

Find the area under the standard normal curve to the right of

zequals=1.

0.8413

0.5398

0.1587

0.1397
Answer #1

solution

(A)P( 1< Z < 2)

= P(Z < 2) - P(Z < 1)

Using z table,  

= 0.9772-0.8413  

area=0.1359

(B)

P(z >1 ) =1 - P(z <1 )

Using z table,
= 1 -0.8413

= 0.1587

area = 0.1587